2. In Mendelian segregation, a “dihybrid self” represents a F1 dihybrid that was crossed with itself (AaBb x AaBb). What are the expected F2 phenotypic ratios, if the F1 cross was actually Aa; Bb x Aa; bb? Show a punnet square.
Answer #1

Answer:

Aa;Bb x Aa;bb ---Parents

Ab ab
AB AA; Bb Aa; Bb
Ab AA; bb Aa; bb
aB Aa; Bb aa; Bb
ab Aa; bb aa; bb

A_B_ phenotype = AA;Bb + Aa; Bb = 1+2 = 3

A_bb phenotype = AA; bb + Aa bb = 1+2 = 3

aaB_ phenotype = aaBb = 1

aabb = 1 = 1

The phenotypic ratio = 3:3:1:1